If f(3) = 8 and f(x) = 2 for 3 sxs 7, how small can f(7) possibly be? x Enhanced Feedback
Answer #1

o tom) be a straight line with t) = 2 f(3) = 8 . f (7) ? fcase f(77 - f (3) = 2 7-3 ..467)= 8+8 16